find minimum value of


mod(2z-1) + mod(3z-2)

Grade:

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago
Dear student,
mod(2z-1)=mod(2x-1+i2y)<root(4x2+-4x+4y2)
mod(3z-2)=mod(3x-2+i3y)< root(9x2+4-12x+9y2)
nowfind the minimum value easily....
Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.
Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert
Sagar Singh
B.Tech, IIT Delhi
sdf
7 years ago
please give answer
mycroft holmes
7 years ago
2|z-1/2|+3|z-2/3| = 2[ |z-1/2|+|2/3 – z| ] + |z-2/3]
By triangle inequality we have that the expression
\ge 2 \left|\frac{2}{3} - \frac{1}{2} \right| + \left|z -\frac{2}{3} \right|
\ge \frac{1}{3} + \left|z -\frac{2}{3} \right|
Equality occurs when z = 2/3
Hence min is 2/3